how do u graph x - y = -2
Answer by ewatrrr(24785)

 
Hi
x - y = -2
y = x + 2
x y
0 2 (0,2)
-1 1 (-1,1)
-2 0 (-2,0)
Plot points, connect with the line representing y = x + 2